# Retrieval-Augmented Generation for AI-Generated Content: A Survey

### v5

@@ III-B3 Generator Enhancement
Prompt Engineering
Technologies in prompt engineering [155] that focus on improving the quality of LLMs’ output, such as Prompt compression, Stepback Prompt [156], Active Prompt [157], Chain of Thought Prompt [129], etc., are all applicable to LLM generators in RAG systems. LLMLingua [158] applies a small model to compresses the overall length of the query to accelerate model inference, relieving the negative impact of irrelevant information on the model and alleviating the phenomenon of “Lost in the Middle”[159]. ReMoDiffuse [51] decomposes complex descriptions into anatomical text scripts by using ChatGPT. ASAP [160] incorporates exemplar tuples, consisting of input code, function definitions, analysis results, and corresponding comments, into prompts to yield better results. CEDAR [88] uses a designed prompt template to organize code demonstration, query, and natural language instructions into a prompt. XRICL [145] utilizes COT technology to add translation pairs as an intermediate step in cross linguistic semantic parsing and inference. ACTIVERAG [161] employs the Cognition Nexus mechanism to calibrate the intrinsic cognition of LLMs and applies COT prompt in answer generation. Make-An-Audio [44] is able to use other modalities as input which can provide much richer information for the following process.
Decoding Tuning
Decoding tuning involves enhancing generator control by fine-tuning hyperparameters for increased diversity and constraining the output vocabulary, among other adjustments.
@@
Generator Finetuning
The finetuning of the generator can enhance the model’s ability to have more precise domain knowledge or better fit with the retriever.
@@ III-B4 Result Enhancement
Output Rewrite
Output Rewrite refers to rewriting the content generated by the generator in certain scenarios to meet the needs of downstream tasks.
SARGAM [165] refines outputs in code-related tasks by employing a special Transformer alongside Deletion, Placeholder, and Insertion Classifiers to better align with the real-world code context. Ring [166] obtains diversity results by reranking candidates based on the average of per token log probabilities produced by the generator. CBR-KBQA [54] revises the result by aligning generated relations with those presented in the local neighborhood of the query entity in knowledge graph.
@@ III-B5 RAG Pipeline Enhancement
RAG Pipeline Enhancement refers to optimizing the processes of RAG at the system level in order to achieve better performance results.
Adaptive Retrieval
Some studies on RAG suggest that retrieval doesn’t always enhance the final results. Over-retrieval can lead to resource wastage and potential confusion when the model’s inherent parameterized knowledge suffices for answering relevant questions. Consequently, this chapter will delve into two methods for determining retrieval necessity: rule-based and model-based approaches.
@@
Model-based: Self-RAG [84] uses a trained generator to determine whether to perform a retrieval based on the retrieve token under different user querys. Ren et al. [171] used ”Judgment Prompting” to determine whether LLMs can answer relevant questions and whether their answers are correct or not, thereby assisting in determining the necessity of a retrieval. SKR [172] uses the ability of LLMs themselves to judge in advance whether they can answer the question, and if they can answer, no retrieval is performed. Rowen [173] translates a question into multiple languages and checks for answer consistency across these languages, using the results to determine the need for information retrieval. AdaptiveRAG [174] dynamically decides whether to retrieve based on the query complexity by a classifier, which is a smaller LM.
Iterative RAG
Iterative RAG progressively refines results by repeatedly cycling through retrieval and generation phases, rather than a single round.
RepoCoder [175] uses an iterative retrieval-generation approach for code completion, refining queries with previously generated code to better utilize dispersed information and improve outcomes. ITER-RETGEN [176] iteratively enhances content quality by using the generator’s output to pinpoint knowledge gaps, retrieving necessary information, and informing future generation cycles. SelfMemory [177] utilizes a retrieval-augmented generator iteratively to form an expansive memory pool, from which a memory selector picks an output to inform the next generation cycle. RAT [178] initial generates content by an LLM with a zero-shot CoT prompt, then revises each thought step by retrieving knowledge from external knowledge base.
@@ IV Applications
RAG for Text
Question Answering Human-Machine Conversation Neural Machine Translation Summarization Others
REALM§ TKEGEN§ TOG SKR§ Self-RAG§ RIAG FiD§ RETRO§ NPM§ CREA-ICL BlenderBot3§ CEG Internet-Augmented-DG§ ConceptFlow§ Skeleton-to-Response§ NMT-with-Monolingual-TM§ TRIME§ KNN-MT§ COG RAMKG§ RPRR RIGHT§ Unlimiformer§ CONCRETE§ Atlas§ KG-BART§ R-GQA§
RAG for Code
Code Generation Code Summarization Code Completion Automatic Program Repair Text-to-SQL and Code -based Semantic Parsing Others
SKCODER§ RRGCode ARKS RECODE KNN-TRANX Toolcoder§ RACE BASHEXPLAINER READSUM Rencos CoRec Tram§ EDITSUM ReACC RepoCoder§ De-Hallucinator REPOFUSE§ RepoFusion§ EDITAS§ RING CEDAR§ RAP-Gen§ InferFix§ SARGAM§ RTLFixer§ XRICL§ SYNCHROMESH§ RESDSQL§ REFSQL§ CodeICL§ MURRE De-fine Code4UIE§ E&V StackSpotAI§ ImputBlaster
RAG for Knowledge RAG for 3D
Knowledge Base QA Knowledge-augmented Open-domain QA Table for QA Others Text-to-3D
CBR-KBQA§ TIARA§ Keqing§ RNG-KBQA ReTraCk§ SKP§ UniK-QA KG-FiD GRAPE SKURG KnowledGPT EFSUM§ EfficientQA CORE§ Convinse RINK§ T-RAG§ StructGPT GRetriever§ SURGE§ K-LaMP RHO ReMoDiffuse AMD
RAG for Image RAG for Video
Image Generation Image Captioning Others Video Captioning Video QA & Dialogue Others
RetrieveGAN IC-GAN§ Re-imagen§ RDM Retrieve&Fuse§ KNN-Diffusion MA REVEAL SMALLCAP CRSR RA-Transformer PICa Maira KIF RA-VQA KaVD§ R-ConvED§ CARE§ EgoInstructor§ MA-DRNN R2A Tvqa+§ VGNMN VidIL RAG-Driver Animate-A-Story§
RAG for Science RAG for Audio
Drug Discovery Biomedical Informatics Enhancement Math Applications Audio Generation Audio Captioning
RetMol§ PromptDiff PoET Chat-Orthopedist§ BIOREADER MedWriter QARAG LeanDojo RAG-for-math-QA Re-AudioLDM§ Make-An-Audio§ RECAP§
@@ IV-A1 Question Answering
Question Answering involves the process of providing responses to posed questions by drawing from a vast and comprehensive collection of textual sources.
FiD [35] and REALM [33] identify the top-k most pertinent article snippets based on the query and forward each snippet along with the question to LLMs to generate k responses. These responses are then synthesized into a final answer. Toutanova et al. [179] substituted the text corpus in REALM with subgraphs from a knowledge graph, yielding impressive results. RETRO [36] employs attention mechanisms to integrate the question with relevant retrieved documents within the model to produce the final answer. SKR [172] observes that using RAG does not invariably benefit Question Answering and thus explored guiding the model to evaluate its grasp of pertinent knowledge, subsequently adapting its use of external resources for retrieval enhancement. TOG [180] introduces an innovative knowledge graph-augmented LLM framework, which excels by fostering interactions between LLMs and the Knowledge Graph and by expanding the inference path space with beam search. NPM [118] pioneers the use of nonparametric data distributions in lieu of the softmax layer, enabling models with fewer parameters to perform effectively. Self-RAG [84] improves answer quality by learning to discern when to retrieve, assess the retrieved content’s relevance, and evaluate the final generated results using four types of reflective tokens. CL-ReLKT [181] employs a language-generalized encoder to bridge the gap between question-document pairs across languages, thus better leveraging multilingual data. CORE [182] mitigates language resource disparities by introducing a novel dense passage retrieval algorithm and a multilingual autoregressive generation model. Lastly, EAE [112] enhances answer quality by retrieving entity embeddings for query entities and integrating these with hidden states for further processing.
@@ IV-A2 Fact Verification
Fact Verification typically refers to determining whether a given natural language text and a related claim or assertion match the facts in the text. CONCRETE [187] leverages cross-lingual retrieval mechanisms to tap into a wealth of multilingual evidence, effectively bridging the gap in resources for languages that are underrepresented in fact-checking datasets. Hagström et al. [188] proved on LLaMA [4] and Atlas [30] that search augmentation is more beneficial for solving inconsistency problems than increasing model size. Atlas [30] shows that using RAG to support LLMs in knowledge-intensive tasks markedly improves their few-shot learning performance.
@@ IV-A3 Commonsense Reasoning
Commonsense Reasoning entails the capability of machines to infer or make decisions on problems or tasks in a human-like manner, drawing upon their acquired external knowledge and its application. KG-BART [189] expands the conceptual landscape by incorporating intricate interrelations among diverse concepts within a knowledge graph. It employs graph attention mechanisms to aid LLMs in crafting more nuanced and logically coherent sentences. Wan et al. [190] constructed the CONFLICTINGQA dataset with contentious questions and conflicting answers to study how textual features affect LMs’ handling of controversial issues.
@@ IV-A4 Human-Machine Conversation
Human-Machine Conversation encompasses the ability of machines to comprehend natural language and adeptly employ this skill to engage with humans seamlessly. ConceptFlow [191] leverages a commonsense knowledge graph to structure conversations, directing the flow of dialogue based on attention scores, and propelling the conversation forward. Cai et al. [192] reimagined the text generation task as a cloze test by retrieving and distilling the essence of past conversational history, leading to notable outcomes. Komeili et al. [193] augmented dialogue generation quality by harnessing advanced search engine technologies to source pertinent content from the internet. BlenderBot3 [194] broadens its search horizon, not only mining relevant internet content but also local dialogue history, and employs entity extraction among other techniques to refine the quality of the resulting dialogue. Kim et al. [195], PARC [196], and CREA-ICL [197] improve the caliber of non-English conversations by incorporating cross-lingual knowledge, effectively addressing the scarcity of non-English datasets and enhancing the quality of the generated dialogue. CEG [198] addresses hallucination issues through a post-processing mechanism, verifying LLM-generated answers through retrieval.
@@ IV-A5 Neural Machine Translation
Neural Machine Translation (NMT) is the automated process of translating text from a source language to a target language [117, 199, 200]. It is a pivotal task in the domain of NLP and represents a significant objective in the pursuit of AI, boasting considerable scientific and practical significance. Cai et al. [199] proposed an innovative approach that utilizes monolingual corpora alongside multilingual learning techniques, challenging the traditional dependency on bilingual corpora in Neural Machine Translation. KNN-MT [200] executes translation tasks at the token level by computing vector space distances. TRIME [117] effectively minimizes the discrepancy between training and inference phases by jointly training the retrieval system and the generation model, thereby enhancing the precision of translations.
@@ IV-A7 Summarization
Summarization is a task aimed at distilling the essential information from lengthy texts and producing a concise, coherent summary that encapsulates the primary themes. There are two main approaches to Summarization: Extractive and Abstractive. Extractive Summarization involves the automatic selection and compilation of key phrases directly from the source text. This technique refrains from creating new sentences, instead repurposing segments from the original text. Abstractive Summarization, on the other hand, entails comprehending the original text’s meaning and reformulating it into new sentences [202, 95, 203, 204]. This approach can convey the source’s intent more fluidly but poses greater challenges in terms of implementation due to its complexity. RAMKG [202] effectively leverages a comprehensive English corpus to bolster the performance of Keyphrase Generation in non-English contexts. Unlimiformer [95] addresses the issue of input length constraints in transformer-based models by retrieving and utilizing the top-k most relevant hidden states, thereby extending the model’s capacity to handle longer inputs. RPRR [203] employs a Retrieve-Plan-Retrieve-Read approach to overcome the limited context window constraints faced by LLMs, utilizing retrieved information to generate high-quality Wikipedia documents for emerging events. RIGHT [204] chooses to use different types of retrievers in different datasets to enhance the generator.
@@ IV-B1 Code Generation
Logit-based RAG is also applicable for code generation. RECODE [78] retrieves NL descriptions and paired codes using edit distance, then extracts n-gram action subtrees from ASTs. During LSTM-based generation, the processed subtrees are leveraged through logits at each decoding step. kNN-TRANX [119] uses a seq2tree model to convert NL to code AST. At each decoding step of the generation, the hidden states are searched within the AST prefix datastore to form new probabilities, which are later combined with the seq2tree model’s output through a confidence network.
@@ IV-B2 Code Summarization
Re2Com [100] and EditSum [97] retrieve similar codes using BM25 and generate summary using LSTM. They separately encode the input, the retrieved code, and the corresponding summary, then combine the hidden states or logits in the decoder. HGNN [217] instead uses code edit distance for retrieval, and substitutes the code encoder with hybrid GNN on their Code Property Graphs [218]. RACE [101] employs separate encoders for the input code difference, the retrieved code differences through dense retrieval, and corresponding commit message to generate the final commit messages. BASHEXPLAINER [98] applies dense retrieval, and fuses the embeddings for subsequent transformer-based decoder. READSUM [219] uses Levenshtein distance for retrieval, and employs a fusion network to combine the representations of retrieved codes and summaries.

### v6

@@ III-B3 Generator Enhancement
Prompt Engineering: Technologies in prompt engineering [166] that focus on improving the quality of LLMs’ output, such as prompt compression, Stepback Prompt [167], Active Prompt [168], Chain of Thought Prompt [133], etc., are all applicable to LLM generators in RAG systems.
LLMLingua [169] applies a small model to compresses the overall length of the query to accelerate model inference, relieving the negative impact of irrelevant information on the model and alleviating the phenomenon of “Lost in the Middle”[170]. ReMoDiffuse [51] decomposes complex descriptions into anatomical text scripts by using ChatGPT. ASAP [171] incorporates exemplar tuples, consisting of input code, function definitions, analysis results, and corresponding comments, into prompts to yield better results. CEDAR [89] uses a designed prompt template to organize code demonstration, query, and natural language instructions into a prompt. XRICL [153] utilizes COT technology to add translation pairs as an intermediate step in cross linguistic semantic parsing and inference. ACTIVERAG [172] employs the Cognition Nexus mechanism to calibrate the intrinsic cognition of LLMs and applies COT prompt in answer generation. Make-An-Audio [44] is able to use other modalities as input which can provide much richer information for the following process.
Decoding Tuning: Decoding tuning involves enhancing generator control by fine-tuning hyperparameters for increased diversity and constraining the output vocabulary, among other adjustments.
@@
Generator Finetuning: The finetuning of the generator can enhance the model’s ability to have more precise domain knowledge or better fit with the retriever.
@@ III-B4 Result Enhancement
Output Rewrite: Output Rewrite refers to rewriting the content generated by the generator in certain scenarios to meet the needs of downstream tasks. SARGAM [176] refines outputs in code-related tasks by employing a special Transformer alongside Deletion, Placeholder, and Insertion Classifiers to better align with the real-world code context. Ring [177] obtains diversity results by reranking candidates based on the average of per token log probabilities produced by the generator. CBR-KBQA [54] revises the result by aligning generated relations with those presented in the local neighborhood of the query entity in knowledge graph.
@@ III-B5 RAG Pipeline Enhancement
RAG pipeline enhancement refers to optimizing the overall process of RAG in order to achieve better performance results.
Adaptive Retrieval: Some studies on RAG suggest that retrieval doesn’t always enhance the final results. Over-retrieval can lead to resource wastage and potential confusion when the model’s inherent parameterized knowledge suffices for answering relevant questions. Consequently, this chapter will delve into two methods for determining retrieval necessity: rule-based and model-based approaches.
@@
Model-based: Self-RAG [85] uses a trained generator to determine whether to perform a retrieval based on the retrieve token under different user querys. Ren et al. [182] used “Judgment Prompting” to determine whether LLMs can answer relevant questions and whether their answers are correct or not, thereby assisting in determining the necessity of a retrieval. SKR [183] uses the ability of LLMs themselves to judge in advance whether they can answer the question, and if they can answer, no retrieval is performed. Rowen [184] translates a question into multiple languages and checks for answer consistency across these languages, using the results to determine the need for information retrieval. AdaptiveRAG [185] dynamically decides whether to retrieve based on the query complexity by a classifier, which is a smaller LM.
Iterative RAG: Iterative RAG progressively refines results by repeatedly cycling through retrieval and generation phases, rather than a single round.
RepoCoder [186] uses an iterative retrieval-generation approach for code completion, refining queries with previously generated code to better utilize dispersed information and improve outcomes. ITER-RETGEN [187] iteratively enhances content quality by using the generator’s output to pinpoint knowledge gaps, retrieving necessary information, and informing future generation cycles. SelfMemory [188] utilizes a retrieval-augmented generator iteratively to form an expansive memory pool, from which a memory selector picks an output to inform the next generation cycle. RAT [189] initially generates content by an LLM with a zero-shot CoT prompt, then revises each thought step by retrieving knowledge from external knowledge base.
@@ IV Applications
RAG for Text Question Answering Human-Machine Conversation Neural Machine Translation Summarization Others REALM§ TKEGEN§ RIAG Fid§ RETRO§ NPM§ SKR§ Self-RAG§ TOG ConceptFlow§ Skeleton-to-Response§ CREA-ICL Internet-Augmented-DG§ BlenderBot3§ CEG NMT-with-Monolingual-TM§ KNN-MT§ COG TRIME§ RAMKG§ Unlimiformer§ RPRR RIGHT§ CONCRETE§ Atlas§ KG-BART§ R-GQA§
RAG for Code Code Generation Code Summary Code Completion Automatic Program Repair Text-to-SQL and Code-based Semantic Parsing Others SKCODER§ RRGCode ARKS KNN-TRANX RECODE Toolcoder§ RACE BASHEXPLAINER READSUM Rencos CoRec Tram§ EDITSUM ReACC RepoCoder§ De-Hallucinator REPOFUSE§ RepoFusion§ EDITAS§ RING CEDAR§ RAP-Gen§ InferFix§ SARGAM§ RTLFixer§ XRICL§ SYNCHROMESH§ RESDSQL§ REFSQL§ CodeICL§ MURRE StackSpotAI§ E&V Code4UIE§ De-fine ImputBlaster
RAG for Knowledge Knowledge Base QA Knowledge-augmented Open-domain QA Table for QA Others CBR-KBQA§ TIARA§ Keqing§ RNG-KBQA ReTraCk§ SKP§ UniK-QA KG-FiD GRAPE SKURG KnowledGPT EFSUM§ EfficientQA CORE§ Convinse RINK§ T-RAG§ StructGPT GRetriever§ SURGE§ K-LaMP RHO RAG for 3D Text-to-3D ReMoDiffuse AMD
RAG for Image Image Generation Image Captioning Others RetrieveGAN IC-GAN§ Re-imagen§ RDM Retrieve&Fuse§ KNN-Diffusion MA REVEAL SMALLCAP CRSR RA-Transformer PICa Maira KIF RA-VQA RAG for Video Video Captioning Video QA&Dialogue Others KaVD§ R-ConvED§ CARE§ EgoInstructor§ MA-DRNN R2A Tvqa+§ VGNMN VidIL RAG-Driver Animate-A-Story§
RAG for Science Drug Discovery Biomedical Informatics Enhancement Math Applications RetMol§ PromptDiff PoET Chat-Orthopedist§ BIOREADER MedWriter QARAG LeanDojo RAG-for-math-QA RAG for Audio Audio Generation Audio Captioning Re-AudioLDM§ Make-An-Audio§ RECAP§
@@ IV-A1 Question Answering
Question answering involves the process of providing responses to posed questions by drawing from a vast and comprehensive collection of textual sources. FiD [35] and REALM [33] identify the top-k most pertinent article snippets based on the query and forward each snippet along with the question to LLMs to generate k responses. These responses are then synthesized into a final answer. Toutanova et al. [190] substituted the text corpus in REALM with subgraphs from a knowledge graph, yielding impressive results. As shown in Fig. 5, RETRO [36] employs attention mechanisms to integrate the question with relevant retrieved documents within the model to produce the final answer. SKR [183] observes that using RAG does not invariably benefit question answering and thus explored guiding the model to evaluate its grasp of pertinent knowledge, subsequently adapting its use of external resources for retrieval enhancement. TOG [191] introduces an innovative knowledge graph-augmented LLM framework, which excels by fostering interactions between LLMs and the knowledge graph and by expanding the inference path space with beam search. NPM [119] pioneers the use of nonparametric data distributions in lieu of the softmax layer, enabling models with fewer parameters to perform effectively. CL-ReLKT [192] employs a language-generalized encoder to bridge the gap between question-document pairs across languages, thus better leveraging multilingual data. CORE [193] mitigates language resource disparities by introducing a novel dense passage retrieval algorithm and a multilingual autoregressive generation model. Lastly, EAE [113] enhances answer quality by retrieving entity embeddings for query entities and integrating these with hidden states for further processing.
@@ IV-A2 Fact Verification
Fact verification typically refers to determining whether a given natural language text and a related claim or assertion match the facts in the text.
CONCRETE [198] leverages cross-lingual retrieval mechanisms to tap into a wealth of multilingual evidence, effectively bridging the gap in resources for languages that are underrepresented in fact-checking datasets. Atlas [30] shows that using RAG to support LLMs in knowledge-intensive tasks markedly improves their few-shot learning performance. Hagström et al. [199] proved on LLaMA [4] and Atlas [30] that search augmentation is more beneficial for solving inconsistency problems than increasing model size. Stochastic RAG [200] employs stochastic sampling without replacement to address the non-differentiable topk selection process in RAG retrieval, enabling end-to-end optimization and achieving excellent results in fact verification scenarios.
@@ IV-A3 Commonsense Reasoning
Commonsense reasoning entails the capability of machines to infer or make decisions on problems or tasks in a human-like manner, drawing upon their acquired external knowledge and its application.
KG-BART [201] expands the conceptual landscape by incorporating intricate interrelations among diverse concepts within a knowledge graph. It employs graph attention mechanisms to aid LLMs in crafting more nuanced and logically coherent sentences. Wan et al. [202] constructed the CONFLICTINGQA dataset with contentious questions and conflicting answers to study how textual features affect LMs’ handling of controversial issues.
@@ IV-A4 Human-Machine Conversation
Human-machine conversation encompasses the ability of machines to comprehend natural language and adeptly employ this skill to engage with humans seamlessly.
ConceptFlow [203] leverages a commonsense knowledge graph to structure conversations, directing the flow of dialogue based on attention scores, and propelling the conversation forward. Cai et al. [204] reimagined the text generation task as a cloze test by retrieving and distilling the essence of past conversational history, leading to notable outcomes. Komeili et al. [205] augmented dialogue generation quality by harnessing advanced search engine technologies to source pertinent content from the internet. BlenderBot3 [206] broadens its search horizon, not only mining relevant internet content but also local dialogue history, and employs entity extraction among other techniques to refine the quality of the resulting dialogue. Kim et al. [207], PARC [208], and CREA-ICL [209] improve the caliber of non-English conversations by incorporating cross-lingual knowledge, effectively addressing the scarcity of non-English datasets and enhancing the quality of the generated dialogue. CEG [210] addresses hallucination issues through a post-processing mechanism, verifying LLM-generated answers through retrieval.
@@ IV-A5 Neural Machine Translation
Neural Machine Translation (NMT) is the automated process of translating text from a source language to a target language [118, 211, 212]. It is a pivotal task in the domain of NLP and represents a significant objective in the pursuit of AI, boasting considerable scientific and practical significance.
Cai et al. [211] proposed an innovative approach that utilizes monolingual corpora alongside multilingual learning techniques, challenging the traditional dependency on bilingual corpora in Neural Machine Translation. kNN-MT [212] executes translation tasks at the token level by computing vector space distances. TRIME [118] effectively minimizes the discrepancy between training and inference phases by jointly training the retrieval system and the generation model, thereby enhancing the precision of translations.
@@ IV-A7 Summarization
Summarization is a task aimed at distilling the essential information from lengthy texts and producing a concise, coherent summary that encapsulates the primary themes. There are two main approaches to summarization: extractive and abstractive.
Extractive summarization involves the automatic selection and compilation of key phrases directly from the source text, which refrains from creating new sentences, instead repurposing segments from the original text.
Abstractive summarization, on the other hand, entails comprehending the original text’s meaning and reformulating it into new sentences [214, 96, 215, 216], which can convey the source’s intent more fluidly but poses greater challenges in terms of implementation due to its complexity. RAMKG [214] effectively leverages a comprehensive English corpus to bolster the performance of keyphrase generation in non-English contexts. Unlimiformer [96] addresses the issue of input length constraints in transformer-based models by retrieving and utilizing the top-k most relevant hidden states, thereby extending the model’s capacity to handle longer inputs. RPRR [215] employs a Retrieve-Plan-Retrieve-Read approach to overcome the limited context window constraints faced by LLMs, utilizing retrieved information to generate high-quality Wikipedia documents for emerging events. RIGHT [216] chooses to use different types of retrievers in different datasets to enhance the generator. M-RAG [217] significantly enhances text summarization by segmenting documents into various databases and incorporating multi-agent reinforcement learning techniques.
@@ IV-B1 Code Generation
Logit-based RAG is also applicable for code generation. RECODE [79] retrieves NL descriptions and paired codes using edit distance, then extracts n-gram action subtrees from ASTs. During LSTM-based generation, the processed subtrees are leveraged through logits at each decoding step. kNN-TRANX [120] uses a seq2tree model to convert NL to code AST. During each decoding step, hidden states are searched in the AST prefix datastore to create new probabilities, later merged with the seq2tree model’s output via a confidence network.
@@ IV-B2 Code Summarization
Re2Com [101] and EditSum [98] retrieve similar codes using BM25 and generate summary using LSTM. They separately encode the input, the retrieved code, and the corresponding summary, then combine the hidden states or logits in the decoder. HGNN [230] instead uses code edit distance for retrieval, and substitutes the code encoder with hybrid GNN on their Code Property Graphs (CPG) [231]. RACE [102] employs separate encoders for the input code difference, the retrieved code differences through dense retrieval, and corresponding commit message to generate the final commit messages. BASHEXPLAINER [99] applies dense retrieval, and fuses the embeddings for subsequent transformer-based decoder. READSUM [232] uses Levenshtein distance for retrieval, and employs a fusion network to combine the representations of retrieved codes and summaries.